# JFR: A Jump Frontier Relaxation Strategy for Fast Bellman–Ford

### v3

@@ Abstract
Across diverse graph families, JFR reduces relaxation operations by 25–99%. Its complexity is $O(n+m\,\bar{D}/k)$, where the overall speedup depends on a cost-benefit balance between the $1/k$ relaxation reduction and the accumulated $k$-LMH overhead.
@@ 4.4 C++ Experiments: Large-Scale Randomized Benchmarking
Table 2: C++ Experiments (default): JFR vs SPFA-SLF (Runtime and Relaxation Ops, Averaged over Large-Scale Tests)
Graph Algorithm Time (ms) Ops Check
Sparse_XL JFR 85.54 121,626.53 PASS
SPFA-SLF 43.32 1,464,074.94 PASS
NegDense_XL JFR 432.35 10,543,741.68 PASS
SPFA-SLF 409.39 20,668,108.21 PASS
Windmill_XL JFR 10.04 185,487.19 PASS
SPFA-SLF 4.99 249,528.29 PASS
SLF_Killer_XL JFR 249.34 861,774.90 PASS
SPFA-SLF 4,521.67 154,592,700.00 PASS
@@
Sparse_XL JFR 16.91 114,054 PASS
SPFA-SLF 11.18 1,490,188 PASS
NegDense_XL JFR 52.80 10,241,868 PASS
SPFA-SLF 56.63 20,721,610 PASS
Windmill_XL JFR 0.45 145,395 PASS
SPFA-SLF 0.28 195,611 PASS
SLF_Killer_XL JFR 12.51 1,014,790 PASS
SPFA-SLF 1,030.46 160,003,344 PASS
Impact of Compilation Environment
Comparing Table 2 (Default) and Table 3 (-O3 Optimized) reveals that JFR benefits disproportionately from compiler optimizations. In the default environment, the overhead of the std::priority_queue masks JFR’s algorithmic efficiency. However, under -O3 optimization, JFR not only drastically widens the performance gap on adversarial graphs (achieving an $\approx$82x speedup on SLF_Killer_XL) but also surpasses SPFA-SLF on dense negative graphs (NegDense_XL). This confirms that in realistic high-performance scenarios, the reduction in relaxation operations effectively amortizes the logarithmic queue overhead.
@@ 6 Conclusion and Future Work
Distinctive Strength of JFR: Across all evaluated graph topologies, JFR consistently achieves a substantial reduction in relaxation operations relative to SPFA-SLF, often by multiple orders of magnitude on dense or adversarial graphs. This operational efficiency directly translates into faster runtimes and reduced memory-access-induced energy usage. Importantly, even in occasional edge cases where JFR performs more relaxations than SPFA-SLF, the overall trend demonstrates robust suppression of redundant operations, showcasing its resilience and applicability for large-scale, complex networks.
@@
    Operational Efficiency: By reducing effective relaxation operations by 25–99%, JFR lowers the combinatorial workload. This advantage is particularly notable in adversarial or worst-case topologies, where label-correcting algorithms approach their theoretical $O(VE)$ complexity.

### v4

@@ Abstract
Across diverse graph families, JFR reduces relaxation operations by -31–99%. Its complexity is $O(n+m\,\bar{D}/k)$, where the overall speedup depends on a cost-benefit balance between the $1/k$ relaxation reduction and the accumulated $k$-LMH overhead.
@@ 4.4 C++ Experiments: Large-Scale Randomized Benchmarking
Theorem 2.5 provides the performance lower bound (Lower Bound) of the JFR strategy under the ideal fixed depth condition. Our solver approximates and dynamically optimizes this $k$ value through the greedy strategy of the priority queue, achieving an engineering-level optimization.
@@
Sparse_XL JFR 26.80 114,054 PASS
SPFA-SLF 18.79 181,717 PASS
NegDense_XL JFR 62.80 10,204,376 PASS
SPFA-SLF 68.63 10,738,464 PASS
Windmill_XL JFR 0.55 143,750 PASS
SPFA-SLF 0.35 109,091 PASS
SLF_Killer_XL JFR 13.56 1,007,091 PASS
SPFA-SLF 1,064.71 44,693,930 PASS
Based on the large-scale benchmark results presented in Table 2, several key physical characteristics of the JFR framework emerge from the comparison with the SPFA-SLF baseline.
4.5 Operational Suppression in Adversarial Topologies
The most prominent observation is the drastic reduction of relaxation operations in the SLF_Killer_XL dataset. While SPFA-SLF executes approximately 44.69 million relaxations, JFR restricts the total count to just 1.01 million. This 97.7% reduction in operations corresponds to a massive reduction in runtime from 1,064.71 ms down to 13.56 ms.
This data demonstrates that JFR acts as a structural stabilizer. In topologies designed to induce exponential oscillation in label-correcting algorithms, JFR maintains a near-linear operational scale, effectively preventing the ”computational explosion” observed in the baseline.
4.6 Consistency of Operational Count
A critical observation across the four test categories is the stability of JFR’s relaxation counts. Across Sparse_XL, Windmill_XL, and SLF_Killer_XL, JFR’s operations stay within a relatively narrow range (approximately $1.1\times 10^{5}$ to $1.0\times 10^{6}$), whereas SPFA-SLF fluctuates wildly from $1.0\times 10^{5}$ to over $4.4\times 10^{7}$.
Even in the most complex NegDense_XL scenario, JFR’s operation count ($1.02\times 10^{7}$) remains slightly lower than that of SPFA-SLF ($1.07\times 10^{7}$). This stability indicates that JFR provides highly predictable performance, ensuring that the algorithm’s workload is governed by the graph’s fundamental reachability rather than its specific edge-weight distribution.
4.7 Trade-offs in Low-Complexity Regimes
The results for Sparse_XL and Windmill_XL reveal the inherent constant-factor overhead of the JFR framework. In these instances, JFR exhibits higher runtimes (26.80 ms and 0.55 ms) compared to SPFA-SLF (18.79 ms and 0.35 ms), despite JFR achieving a 38.16% reduction in operations for the sparse case.
This confirms that the priority-queue maintenance and frontier-filtering logic introduce a fixed computational cost. However, the data shows this trade-off is asymmetric: the marginal time penalty in simple graphs is negligible compared to the magnitude of time savings achieved in dense or adversarial environments.
@@ 6 Conclusion and Future Work
Distinctive Strength of JFR: JFR demonstrates a pervasive reduction in relaxation workload over SPFA-SLF across a wide range of evaluated topologies. This advantage is particularly pronounced in dense and adversarial scenarios, where the framework curtails redundant relaxations by multiple orders of magnitude.
@@
    Operational Efficiency: As demonstrated in Table 2, the reduction in relaxation operations is highly dependent on the graph topology. While the percentage reduction may be limited in sparse or dense regimes—and a proportional increase is observed in structured environments like Windmill_XL—the absolute reduction in the total number of operations remains significant as the scale of the graph ($N$ and $M$) increases.
    This characteristic proves that JFR is particularly advantageous for large-scale networks, where even a marginal percentage gain translates into the elimination of millions of redundant relaxations. By suppressing the combinatorial workload, JFR effectively prevents the algorithms from approaching their theoretical $O(VE)$ worst-case complexity in adversarial or large-scale real-world scenarios.